# AI Governance through Markets

### v1

@@ Abstract.
This paper argues that market governance mechanisms should be considered a key approach in the governance of artificial intelligence (AI), alongside traditional regulatory frameworks. While current governance approaches have predominantly focused on regulation, we contend that market-based mechanisms offer effective incentives for responsible AI development. We examine four emerging vectors of market governance: insurance, auditing, procurement, and due diligence, demonstrating how these mechanisms can affirm the relationship between AI risk and financial risk while addressing capital allocation inefficiencies. While we do not claim that market forces alone can adequately protect societal interests, we maintain that standardised AI disclosures and market mechanisms can create powerful incentives for safe and responsible AI development. This paper urges regulators, economists, and machine learning researchers to investigate and implement market-based approaches to AI governance.
@@ 1. Overview
The field of Artificial Intelligence (AI) governance has predominantly emphasised regulatory frameworks (Anderljung et al., 2023; Schuett, 2023; Roberts et al., 2021) and international cooperation (Cha, 2024; Gruetzemacher et al., 2023) to address AI risk. Meanwhile, uncertainty around AI risks is a major barrier to widespread enterprise adoption (KPMG, [n. d.]) (Altintas, 2024). Despite economic benefits, organizations recognize AI risk as a business risk and lack the tools to confidently address it (Deloitte, 2024a). Market governance approaches, such as insurance, auditing, procurement, and due diligence, can serve to both mitigate AI risk and enable AI growth - aligning market forces with prosocial behaviour (Smith, 2016).
@@
The analysis of the market governance of AI opens up distinct opportunities for both public and private involvement. Mechanisms such as insurance, auditing, due diligence, and procurement offer opportunities for both startups and large enterprises to capitalise on the growing need for AI de-risking which is projected to reach a value of \$276 billion by 2030 (Ventures, 2024). These mechanisms also afford policymakers and quasi-regulatory entities—including industry consortia, trade associations, and standards organisations — strategic pathways for market shaping. Through the deployment of incentive structures, subsidisation programs, public-private collaborations, and standardisation frameworks, policymakers can leverage markets to govern artificial intelligence development while advancing critical economic, technological, and societal imperatives.
@@ 1.1. Governance Through Markets
As Hadfield and Clark argue (Hadfield and Clark, 2023) AI fundamentally disrupts traditional regulatory frameworks in ways that demand comprehensive reconsideration of our approach. Their analysis suggests that the primary challenge isn’t simply creating new rules for AI-specific risks, but developing dynamic regulatory approaches that can adapt to AI’s pervasive influence. Their proposal for regulatory markets can be seen as one possible structure within which market governance mechanisms can operate. Their example demonstrates the broad array of policy strategies available towards market governance of AI - these can range from complete self-governance to a government mandates with punitive enforcement.
Market governance mechanisms often function as private regulators, for-profit and non- profit organizations that develop and supply regulatory services, which compete to sell to targets, this is the case for instance with the protocol organisations and assurance providers (Micheler and Whaley, 2020) we describe later. This is not necessarily always the case, due diligence and insurance have similar effects but would could not fall under this label.
It is important to note, that AI Risk and its associated uncertainty is a major barrier to AI adoption across the economy (Ameye et al., 2023; McKinsey and Company, 2024). Exemplified by how AI service providers have resorted to providing unlimited indemnities in order to maintain sales volume (Cloud, [n. d.]). Market governance mechanisms work to reduce asymmetry and risk, enhancing the stability and efficiency of markets and enabling adoption (Kauffman and Weber, 2021). To give an example, insurance functions as market governance by facilitating the distribution and management of uncertainty across a broad pool of market participants (Knight, 1921), thereby mitigating individual exposure to risk while promoting resource allocation through the collective absorption of losses (Boettke and Candela, 2021).
@@ 1.2. AI Risk as Market Failure
The uncertainty around AI risk is a major barrier to enterprise adoption of AI technologies (KPMG, [n. d.]; Altintas, 2024). Large enterprise customers are uncertain about how to measure, mitigate and govern the risks posed by AI integration into their organisations (Deloitte, 2024a). Notwithstanding the potential economic advantages, this issue persists as a significant impediment to the dissemination of AI technologies. Consequently, organizations have already recognized that AI-related risks are synonymous with business risks and are responding in a manner reflective of this understanding. The awareness of these risks and the inability of organisations to address them prevents firms and society at large from benefiting from AI technologies.
In this vein, we may understand the uncertainty around AI Risk and its mitigation as market failure. The failure, that is, to measure and mitigate the private costs of AI development and deployment and price its externalities. By AI risk, we imply risk from the development and deployment of AI systems. This risk to the firm is magnified in how closely AI is being integrated to value-generating and infrastructural operations. We use the following formula to quantify this risk-adjusted value (Artzner et al., 2007) of AI investments:
@@
  • $\lambda$ (Risk Aversion Coefficient): A factor representing a market actor’s sensitivity to risk, where higher values denote a stronger preference for certainty and a greater aversion to risk. It scales the impact of risk (as measured by $\sigma(X)$) on the risk-adjusted value, personalizing the adjustment based on individual or organizational risk preferences.
@@
Behavioural factors also play a crucial role in the accurate assessment of AI investment risks (Ackert and Deaves, 2009). During hype cycles surrounding AI technologies, cognitive biases may lead firms to underestimate their risk aversion coefficient ($\lambda$) or the variability of outcomes ($\sigma(X)$). This is marked by the financial underperformance of many AI investments to date. Thus, investment decisions are not aligned with the market actor’s true risk tolerance (Rapp et al., 2014; Baker and Ricciardi, 2014).
@@
As seen in Figure 2, at lower values of $\lambda$, the Risk-Adjusted Value (RAV) closely aligns with the expected return $E(X)$, as the risk penalty is relatively small. This reflects scenarios where risk aversion is minimal, and the primary focus is on maximizing returns regardless of the associated volatility. Conversely, at higher values of $\lambda$, the risk term $\lambda\cdot\sigma(X)$ becomes more influential in the RAV calculation, causing RAV to decrease even when $E(X)$ is large. In risk-tolerant scenarios, where $\lambda$ is small, investments or decisions with higher volatility ($\sigma(X)$) may still yield favorable RAVs if $E(X)$ is sufficiently high to offset the modest risk penalty. This interplay highlights the importance of calibrating $\lambda$ to reflect the appropriate level of risk tolerance for a given context.
@@ 2. Risk Distribution: Insurance
Insurance serves as an enabler of transactions amid radical uncertainty by transferring and mitigating risks associated with unforeseen events (Knight, 1921). In the context of deploying AI systems in economic functions, there is an unavoidable material financial risk due to the probabilistic nature of modern AI models (Ghahramani, 2015), as model outputs are undetermined by their inputs. As AI systems become more complex, agentic and autonomous, the potential for unintended consequences increases and so do prospective financial losses.
@@
Insurance can serve as a powerful governance mechanism, shaping responsible AI development and deployment. This can be achieved through risk-based pricing, underwriting standards, and carefully crafted policy exclusions. It also offers potential tools to address thorny liability issues surrounding AI, such as joint causation scenarios or damages caused by hacked AI systems. While traditional insurance domains may exhibit relative stability, AI systems present uniquely complex risk profiles characterized by both substantial magnitude and high variance - characteristics that traditionally catalyse insurance market development (Lior, 2020). The convergence of high-stakes outcomes and inherent unpredictability in AI agent behavior creates natural market conditions for novel insurance products, reflecting the insurance sector’s historical role in facilitating technological adoption through systematic risk distribution.
@@
Insurance has the potential to create a more favourable incentive structure by making practices such as “safety-washing” or underestimating AI-related risks less appealing (Tomei, 2024). The rigour of actuarial assessments may engender an anti-speculative framework for pricing AI development, curtailing so called ’AI Hype’ - much like it has done in other sectors, such as real estate in climate-vulnerable areas (Nyce et al., 2015). By pricing, mitigating, and transferring risk, insurance can encourage more accurate assessment and analysis of AI risk profiles. In so doing, the development of a robust AI insurance market may enable both first-party and third-party stakeholders to more accurately understand the risk profile of AI development.
@@
Cyberrisk provides a crucial precedent to AI risk. It has emerged as the most pressing form of business risk, mirroring the indispensability of software in every major enterprise (Forum, 2024). Accordingly, he financial toll of cyberattacks swells by 15% annually up to \$23tn by 2027 (Economist, 2024). Among numerous potential solutions to this pervasive insecurity, insurance offers some respite. A robust cyber incident insurance market incentivizes organisations to strengthen their cyber defences, ultimately reducing societal cyber risk (Marsh, 2021). While still in its infancy, cyberrisk insurance has promoted the accurate analysis and pricing of the risk landscape. It further incentivises the adoption of cybersecurity measures by making plain the financial consequences incurred and reducing premiums or disqualifying policyholders with insufficient security measures.
@@
As AI becomes integrated to the value creation of major industries, AI insurance will be in increasing demand. Lior et al(Lior, 2021) have emphasised the value of building upon existing insurance infrastructure rather than creating entirely new AI-specific policies, given uncertainties in the field. AI will affect multiple lines of existing insurance, including Technology Errors and Omissions (E&O)/Cyber Liability, Professional Liability, Media Liability, and Employment Practices Liability, among others, depending on the specific use case of the AI (AON, 2024). The manner in which AI risk is incorporated into insurance frameworks will have significant implications for both insurers and policyholders, shaping the management of AI-related risks throughout the private sector.
@@
Societal-scale AI incidents may prompt the development of a CAT-bond market, in order to transfer liability from the insurance sector to capital markets. As systemic AI risk increases with adoption this may provide a key backstop for societal resilience. Catastrophe bonds decrease insurers’ contributions to systemic risk by offloading liability to global capital markets, thus lowering default risks and enhancing stability(Weiß et al., 2013). CAT bonds support faster economic recovery after disasters by providing rapid liquidity and mitigating financial strain on issuers, as strongly supported by empirical studies (Götze and Gürtler, 2021). As the Crowdstrike incident shows parametric CAT-bond triggers misalignment between triggers and actual losses remains a significant limitation leading to basis risk (Finken and Laux, 2009).
@@
Despite these potential benefits, and a number of commercial ventures already underway (Re, 2024), numerous issues still arise in AI risk distribution. The opacity of many AI systems (”black box” characteristics) presents technical challenges in vulnerability assessment, as it makes it difficult to understand and predict how these systems make decisions, which in turn complicates identifying potential vulnerabilities. Furthermore, the widespread integration of shared AI components, particularly foundational model APIs, creates potential cascading failures across industries, complicating the evaluation of systemic risks—a concern particularly acute in high-stakes sectors such as finance and healthcare. Privacy risk assessment faces trade-offs between model accuracy and data protection, compounded by legal ambiguity around PII classification - while current fairness metrics often misalign with legal standards, increasing discrimination liability (Re, 2024). Intellectual property risk is similarly constrained by varying technical and legal interpretations and tort law uncertainty complicates risk assessment, while the lack of legal precedents for AI failures impedes risk estimation. As such, a number of regulatory, standardisation and scientific innovations are still necessary for a robust AI risk market to develop.
@@ 2.1. Case Study: Commercial Real Estate
Commercial real estate (CRE) owners have faced increasing insurance premiums in recent years, prompting behaviour changes to manage these rising costs. Insurance companies have raised their prices to reflect the heightened risks associated with properties in high-risk areas, such as those prone to natural disasters (Greystone, 2024). In response, property owners updated valuations to reflect accurate replacement costs, made significant property improvements (e.g., installing new roofs), and opted for higher or aggregate deductibles to negotiate better rates. These improvements reduced the overall risk of property damage, which in turn lowered the likelihood of claims. This demonstrates how a market mechanism—higher insurance premiums—successfully incentivized risk-reducing behaviours, ultimately benefiting both insurers and property owners through reduced risk (Company, 2024). Additionally, alternative risk transfer strategies, such as structured programs, provided fixed premiums for budget certainty. These actions helped CRE owners reduce premiums and ensure greater financial stability (Greystone, 2024; Company, 2024). A market actor—a company—changed their behaviour to reduce risk and secure better insurance premiums.
@@ 3. Assurance: Auditing and certification
There is significant literature on AI Auditing. The rationale for which varies, some (Ee, 2023; Bernardi et al., 2024) argue that auditing is necessary for public safety, others (Manheim et al., 2024) further point towards keeping AI firms accountable to their own commitments. As Manheim et al note, findings from audits, when disclosed, provide regulators and the scientific community with critical insights into system behaviours and limitations. This transparency supports regulatory initiatives and scientific advancements by deepening the comprehension of system operations. Additionally, it aides due diligence and procurement processes by providing third-party verification to first-party claims.
Further, Audit mechanisms facilitate the resolution of public disclosure trade-offs (see discussion). While system opacity maximises security at the expense of external oversight, comprehensive disclosure enhances scrutiny but risks exposing proprietary technology. Audits potentially offer an intermediary solution, enabling independent risk assessment while maintaining security protocols (Sharkey et al., 2024). Successful audit completion enhances investor confidence and stakeholder trust, aligning institutional credibility with demonstrated accountability measures.

### v2

@@ Abstract.
This paper argues that market governance mechanisms should be considered a key approach in the governance of artificial intelligence (AI), alongside traditional regulatory frameworks. While current governance approaches predominantly focus on regulation, we contend that market-based mechanisms offer effective incentives for responsible AI development. We examine four emerging vectors of market governance: insurance, auditing, procurement, and due diligence, demonstrating how these mechanisms can affirm the relationship between AI risk and financial risk while addressing capital allocation inefficiencies. While we do not claim that market forces alone can adequately protect societal interests, we maintain that standardised AI disclosures and market mechanisms can create powerful incentives for safe and responsible AI development. This paper urges regulators, economists, and machine learning researchers to investigate and implement market-based approaches to AI governance.
@@ 1. Overview
The field of Artificial Intelligence (AI) governance has predominantly emphasised regulatory frameworks (Anderljung et al., 2023; Schuett, 2023; Roberts et al., 2021) and international cooperation (Cha, 2024; Gruetzemacher et al., 2023) to address AI risk. Meanwhile, uncertainty regarding AI risks is a major barrier to widespread enterprise adoption (KPMG, [n. d.]) (Altintas, 2024). Despite the economic benefits, organisations recognise AI risk as a business risk and lack the tools to confidently address it (Deloitte, 2024a). Market governance approaches, such as insurance, auditing, procurement, and due diligence, can serve to both mitigate AI risk and enable AI growth - aligning market forces with prosocial behaviour (Smith, 2016).
@@
Analysis of the market governance of AI opens up distinct opportunities for both public and private involvement. Mechanisms such as insurance, auditing, due diligence, and procurement offer opportunities for both startups and large enterprises to capitalise on the growing need for AI de-risking which is projected to reach a value of \$276 billion by 2030 (Ventures, 2024). These mechanisms also afford policymakers and quasi-regulatory entities—including industry consortia, trade associations, and standards organisations — strategic pathways for market shaping. Through the deployment of incentive structures, subsidisation programs, public-private collaborations, and standardisation frameworks, policymakers can leverage markets to govern artificial intelligence development while advancing critical economic, technological, and societal imperatives.
@@ 1.1. Governance Through Markets
As Hadfield and Clark argue (Hadfield and Clark, 2023) AI fundamentally disrupts traditional regulatory frameworks in ways that demand comprehensive reconsideration of our approach. Their analysis suggests that the primary challenge is not simply creating new rules for AI-specific risks, but developing dynamic regulatory approaches that can adapt to AI’s pervasive influence. Their proposal for regulatory markets can be seen as one possible structure within which market governance mechanisms can operate. Their example demonstrates the broad array of policy strategies available towards market governance of AI, which can range from complete self-governance to government mandates with punitive enforcement.
Market governance mechanisms often function as private regulators, for-profit and non- profit organizations that develop and supply regulatory services, which compete to sell to targets, this is the case for instance with the protocol organisations and assurance providers (Micheler and Whaley, 2020) we describe later. This is not necessarily always the case, due diligence and insurance have similar effects but would not fall under this label.
It is important to note that AI Risk and its associated uncertainty are a major barrier to AI adoption across the economy (Ameye et al., 2023; McKinsey and Company, 2024). This is exemplified by how AI service providers have resorted to providing unlimited indemnities to maintain their sales volume (Cloud, [n. d.]). Market governance mechanisms work to reduce asymmetry and risk, enhancing the stability and efficiency of markets and enabling adoption (Kauffman and Weber, 2021). For example, insurance functions as market governance by facilitating the distribution and management of uncertainty across a broad pool of market participants (Knight, 1921), thereby mitigating individual exposure to risk while promoting resource allocation through the collective absorption of losses (Boettke and Candela, 2021).
@@ 1.2. AI Risk as Market Failure
The uncertainty around AI risk is a major barrier to enterprise adoption of AI technologies (KPMG, [n. d.]; Altintas, 2024). Large enterprise customers are uncertain about how to measure, mitigate and govern the risks posed by AI integration into their organisations (Deloitte, 2024a). Notwithstanding its potential economic advantages, this issue persists as a significant impediment to the dissemination of AI technologies. Consequently, organisations have already recognised that AI-related risks are synonymous with business risks, and are responding in a manner that reflects this understanding. The awareness of these risks and the inability of organisations to address them prevents firms and society at large from benefiting from AI technologies.
In this vein, we may understand the uncertainty around AI Risk and its mitigation as market failure. The failure, that is, to measure and mitigate the private costs of AI development and deployment and price its externalities. By AI risk, we imply a risk from the development and deployment of AI systems. This risk to the firm is magnified in how closely AI is being integrated to value-generating and infrastructural operations. We use the following formula to quantify this risk-adjusted value (Artzner et al., 2007) of AI investments:
@@
  • $\lambda$ (Risk Aversion Coefficient): A factor representing a market actor’s sensitivity to risk, where higher values denote a stronger preference for certainty and a greater aversion to risk. It scales the impact of risk (as measured by $\sigma(X)$) on the risk-adjusted value, personalising the adjustment based on individual or organizational risk preferences.
@@
Behavioural factors also play a crucial role in the accurate assessment of AI investment risks (Ackert and Deaves, 2009). During hype cycles surrounding AI technologies, cognitive biases may lead firms to underestimate their risk-aversion coefficient ($\lambda$) or the variability of outcomes ($\sigma(X)$). This is marked by the financial underperformance of many AI investments to date. Thus, investment decisions are not aligned with the market actor’s true risk tolerance (Rapp et al., 2014; Baker and Ricciardi, 2014).
@@
As seen in Figure 2, at lower values of $\lambda$, the Risk-Adjusted Value (RAV) closely aligns with the expected return $E(X)$, as the risk penalty is relatively small. This reflects scenarios where risk aversion is minimal and the primary focus is on maximising returns, regardless of the associated volatility. Conversely, at higher values of $\lambda$, the risk term $\lambda\cdot\sigma(X)$ becomes more influential in the RAV calculation, causing RAV to decrease even when $E(X)$ is large. In risk-tolerant scenarios, where $\lambda$ is small, investments or decisions with higher volatility ($\sigma(X)$) may still yield favourable RAVs if $E(X)$ is sufficiently high to offset the modest risk penalty. This interplay highlights the importance of calibrating $\lambda$ to reflect the appropriate level of risk tolerance for a given context.
@@ 2. Risk Distribution: Insurance
Insurance serves as an enabler of transactions amid radical uncertainty by transferring and mitigating risks associated with unforeseen events (Knight, 1921). In the context of deploying AI systems in economic functions, there is an unavoidable material financial risk due to the probabilistic nature of modern AI models (Ghahramani, 2015), as model outputs are undetermined by their inputs. As AI systems become more complex, agentic, and autonomous, the potential for unintended consequences increases and so do prospective financial losses.
@@
Insurance can serve as a powerful governance mechanism, shaping responsible AI development and deployment. This can be achieved through risk-based pricing, underwriting standards, and carefully crafted policy exclusions. It also offers potential tools to address thorny liability issues surrounding AI, such as joint causation scenarios or damages caused by hacked AI systems. While traditional insurance domains may exhibit relative stability, AI systems present uniquely complex risk profiles characterised by both substantial magnitude and high variance, which are characteristics that traditionally catalyse insurance market development (Lior, 2020). The convergence of high-stakes outcomes and inherent unpredictability in AI agent behavior creates natural market conditions for novel insurance products, reflecting the insurance sector’s historical role in facilitating technological adoption through systematic risk distribution.
@@
Insurance has the potential to create a more favourable incentive structure by making practices such as “safety-washing” or underestimating AI-related risks less appealing (Tomei, 2024). The rigour of actuarial assessments may engender an anti-speculative framework for pricing AI development, curtailing so called ’AI Hype’ - much like it has done in other sectors, such as real estate in climate-vulnerable areas (Nyce et al., 2015). By pricing, mitigating, and transferring risk, insurance can encourage more accurate assessment and analysis of AI risk profiles. In so doing, the development of a robust AI insurance market may enable both first- and third-party stakeholders to more accurately understand the risk profile of AI development.
@@
Cyberrisk provides a crucial precedent for AI risk. This has emerged as the most pressing form of business risk, mirroring the indispensability of software in every major enterprise (Forum, 2024). Accordingly, he financial toll of cyberattacks swells by 15% annually up to \$23tn by 2027 (Economist, 2024). Among numerous potential solutions to this pervasive insecurity, insurance offers some respite. A robust cyber incident insurance market incentivises organisations to strengthen their cyber defences, ultimately reducing societal cyber risk (Marsh, 2021). While still in its infancy, cyberrisk insurance has promoted accurate analysis and pricing of the risk landscape. It further incentivises the adoption of cybersecurity measures by making plain the financial consequences incurred and reducing premiums or disqualifying policyholders with insufficient security measures.
@@
As AI becomes integrated into the value creation of major industries, AI insurance will be in increasing demand. Lior et al(Lior, 2021) have emphasised the value of building upon existing insurance infrastructure rather than creating entirely new AI-specific policies, given uncertainties in the field. AI will affect multiple lines of existing insurance, including Technology Errors and Omissions (E&O)/Cyber Liability, Professional Liability, Media Liability, and Employment Practices Liability, among others, depending on the specific use case of the AI (AON, 2024). The manner in which AI risk is incorporated into insurance frameworks has significant implications for both insurers and policyholders, shaping the management of AI-related risks throughout the private sector.
@@
Societal-scale AI incidents may prompt the development of a CAT-bond market, in order to transfer liability from the insurance sector to capital markets. As systemic AI risk increases with adoption this may provide a key backstop for societal resilience. Catastrophe bonds decrease insurers’ contributions to systemic risk by offloading liability to global capital markets, thus lowering default risks and enhancing stability(Weiß et al., 2013). CAT bonds support faster economic recovery after disasters by providing rapid liquidity and mitigating financial strain on issuers, as strongly supported by empirical studies (Götze and Gürtler, 2021). As the Crowdstrike incident shows, misalignment between parametric triggers and actual losses remains a significant cause of basis risk (Finken and Laux, 2009).
@@
Despite these potential benefits, and a number of commercial AI Insurance ventures already underway (Re, 2024), numerous issues still arise in AI risk distribution. The opacity of many AI systems (”black box” characteristics) presents severe technical challenges in vulnerability assessment. Furthermore, the widespread integration of shared AI components, particularly foundational model APIs, creates potential cascading failures across industries, complicating the evaluation of systemic risks, a concern that is particularly acute in high-stakes sectors such as finance and healthcare. Privacy risk assessment faces trade-offs between model accuracy and data protection, compounded by legal ambiguity around PII classification - while current fairness metrics often misalign with legal standards, increasing discrimination liability (Re, 2024). Intellectual property risk is similarly constrained by varying technical and legal interpretations, while tort law uncertainty complicates risk assessment. As such, a number of regulatory, legal and scientific innovations are still necessary for a robust AI risk market to develop.
@@ 2.1. Case Study: Commercial Real Estate
Commercial real estate (CRE) owners have faced increasing insurance premiums in recent years, prompting behaviour changes to manage these rising costs. Insurance companies have raised their prices to reflect the heightened risks associated with properties in high-risk areas, such as those prone to natural disasters (Greystone, 2024). In response, property owners updated valuations to reflect accurate replacement costs, made significant property improvements (e.g., installing new roofs), and opted for higher or aggregate deductibles to negotiate better rates. These improvements reduced the overall risk of property damage, which in turn lowered the likelihood of claims. This demonstrates how a market mechanism—higher insurance premiums—successfully incentivised risk-reducing behaviours, ultimately benefiting both insurers and property owners through reduced risk (Company, 2024). Additionally, alternative risk transfer strategies, such as structured programs, provide fixed premiums for budget certainty. These actions helped CRE owners reduce premiums and ensure greater financial stability (Greystone, 2024; Company, 2024). A market actor—a company—changed their behaviour to reduce risk and secure better insurance premiums.
@@ 3. Assurance: Auditing and certification
There is significant literature on AI Auditing. The motivation for Audits for this varies; some (Ee, 2023; Bernardi et al., 2024) argue that auditing is necessary for public safety, while others (Manheim et al., 2024) further point towards keeping AI firms accountable to their own commitments. As Manheim et al. note, findings from audits, when disclosed, provide regulators and the scientific community with critical insights into system behaviours and limitations. This transparency supports regulatory initiatives and scientific advancements by deepening the comprehension of system operations. Additionally, Auditing aids due diligence and procurement processes by providing third-party verification to first-party claims.
Further, audit mechanisms facilitate the resolution of public disclosure trade-offs, while system opacity maximises security at the expense of external oversight, comprehensive disclosure enhances scrutiny but risks exposing proprietary technology. Audits potentially offer an intermediary solution, enabling independent risk assessment while maintaining security protocols (Sharkey et al., 2024). Successful audit completion enhances investor confidence and stakeholder trust, aligning institutional credibility with demonstrated accountability measures.